OpenStudy (asiyahh):

Medal for best answer! < 1 and < 2 are supplementary angles. m < 1 is 4y+7, m < 2 is 9y+4. Find m < 2.

TheSmartOne (thesmartone):

Supplementary angles add up to 180 degress

TheSmartOne (thesmartone):

so 4y+7+9y+4=180 13y+11=180 13y=169 y=13

TheSmartOne (thesmartone):

Now to find angle 2 we substitute y=13 back in to it So 9(13)+4=117+4 This equals to 121

TheSmartOne (thesmartone):

@Asiyahh

OpenStudy (triciaal):

the above is correct. do you have any questions about what @TheSmartOne did?

OpenStudy (asiyahh):

No i got it @triciaal thank you both @TheSmartOne